Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him?
Answer by longjonsilver(2297)

let x be the original price..the cost.

35% of this is 0.35x and he sold the statue for 670, so

x+0.35x = 670
1.35x = 670
x = 670/1.35
x = 496.30 (to 2dp.)

jon.